OpenStudy (diamondanon2):

Find the angle(s) of intersection between the equations f (x) = x 2 + 2x + 1 and g(x) = 1.

OpenStudy (diamondanon2):

Multiple Answers can be selected. 18.4 degrees 26.6 degrees 63.4 degrees 71.6 degrees

OpenStudy (timamits):

63.4 degrees

OpenStudy (diamondanon2):

is that the only answer?

OpenStudy (timamits):

intersection of curves can be found by solving(f(x) = g(x)):- x^2+2x+1 = 1. Which gives x=0,-2. The derivative of f(x) = 2x + 2, which comes 2 and -2 at x=0 and at x= -2 respectively. The other curve has zero slope, so the answer will be arctan(mod(2)) and arctan(mod(-2)), which is nearly 63.4 degrees.

OpenStudy (timamits):

yes

OpenStudy (diamondanon2):

thanks a bunch :)